# Optimization of Excitation Waveforms for Maximum Instantaneous Field Intensity

### v2

@@ III-B Verification
The steps mentioned above are applied to an example of a single dipole antenna. The calculation is simultaneously performed in the Antenna Toolbox for MATLAB [40] (AToM) and CST. The dipole with a length of $15\,\mathrm{mm}$ and a width of $3\,\mathrm{mm}$ is oriented along the $z$ axis. The port is placed in the middle of the dipole and is connected to an ideal $50\,\mathrm{\Omega}$ feed line. Radiation in the spherical coordinate system is studied in the direction $\varphi=0$ and $\theta=\pi/2$. Azimuthal coordinate $\varphi$ references to angle from positive direction of $x$ axis in $xy$ plane and elevation coordinate $\theta$ references to angle from positive direction of $z$ axis.
@@
The antenna, including the port, is given, and all steps visualized in Fig. 1 can be performed. Choosing AToM and CST as the EM solvers and computing the far-field component $F_{\theta}$ in elevation direction for unity excitation at all frequencies results in Fig. 2.
@@
Figure 2: Far field of a dipole antenna with a length of $15\,\mathrm{mm}$ and width of $3\,\mathrm{mm}$ in directions $\varphi=0$ and $\theta=\pi/2$. The antenna is fed through a delta-gap source connected to the feed line, providing a voltage of $1\,\mathrm{V}$ across the entire frequency spectrum.
@@
Since the dipole is aligned along the $z$-axis, only the $\theta$ component of the electric field corresponds to radiation. The curves are in good agreement.
@@
In addition to the far field, an essential known characteristic is total incident energy, represented by the identity operator as seen in (4). This step accomplishes S.1.
@@
Since the frequency data from CST are dense due to the 1001 samples, only linear interpolation is used. Data resulting from AToM are sparse and rational fitting is used for interpolation [^6] Vector fitting [41] could be applied to enforce passivity and causality..
@@
Metrics such as far field and total input power are projected to the band-limited functions defined in Appendix B resulting in matrices as denoted in S.3. The matrix operators allow for optimization S.4 which gives the optimal solution in the frequency domain.
Since the basis functions used for projection are orthogonal and an identity operator represents total input power, the matrix for total input power is an identity matrix. The optimal excitation is a complex conjugate of the far field [1] in Fig. 2. Total input energy is $W_{0}=10^{-10}\,\mathrm{J}$. Meeting the constraint in (8) is ensured by scaling the magnitude of the incident wave. The optimal waveform in the time domain is the result of iFT in S.5, and is plotted in Fig.3.
@@
The last step is the computation of the radiated pulse, shown in Fig. 4.
Figure 4: Optimal radiated waveform for $t_{0}=0\,\mathrm{s}$ at distance $r$ gained from iFT of the far field for AToM data and monitoring the far field during simulation in CST, where the optimal incident wave is applied as excitation.
@@
IV Representative Examples
Multiple antennas and arrays are analyzed using the presented approach. Moreover, this section extends the study to antiferromagnetic memory switching where the target is the electric field magnitude at a chosen point excited by a pulsed incident field. Additional research is also conducted on the energy delivered to the given point during the chosen time interval. Since the solution is independent of the solver, as verified in the previous section, the most suitable one can be used for each scenario.
IV-A Far-Field Pulsed Radiation
IV-A1 Singleport Antennas
As has been concluded in previous work, the bowtie dipole antenna exhibits good performance among planar antennas studied for pulsed radiation in the direction orthogonal to the design plane [26, 28].
In this section, the following antennas are considered: a self-grounded bowtie antenna, a conical dipole antenna, a conical spiral antenna, and a logarithmic periodic antenna. The optimal waveform is determined to maximize the radiated peak power. None of the antennas are topologically optimized for the target. This example aims to demonstrate the utility of the approach and compare the best potentially achievable performance of the proposed antennas, including their optimal excitation waveform.
@@ IV-A1 Singleport Antennas
In [43], the conical spiral antenna is proposed for ground-penetrating radars, an UWB application. The conical spiral radiates a circularly polarized field compared to the other designs which operate in a linearly polarized regime. The antenna proposed in [43] is scaled to fit the circumscribing sphere mentioned above. The simulation is again performed in CST and shown in Fig. 7.
@@
The periodic logarithmic antenna, recognized as an UWB antenna [44], is extensively utilized in electromagnetic emission and immunity testing [6]. Therefore, a wire dipole array is a potential candidate for pulsed radiation and is compared with other antennas. The array is designed to fit the circumscribing sphere and exhibits a decay of $0.9$ in element lengths and an opening angle of $15$ degrees. The strip width equals $0.017$ times the length of the longest element and the number of elements is 20. The antenna is simulated using AToM and shown in Fig. 8.
@@
The question remains as to which antenna performs better. A comparison is shown in Fig. 9.
Figure 9: Normalized optimal radiated pointing vector at chosen distance $r$ of the compared antennas.
The logarithmic periodic antenna exceeds the other designs in performance. The performance of the self-grounded bowtie antenna is about two times lower and the other designs perform even worse. This finding could prompt efforts to develop improved designs.
Regarding the origins of what constitutes good performance, the realized gain in the given direction accounts for the pulse’s sharpness and energy. The emerging differences of the antennas are depicted in Fig. 10 which justifies the performance differences presented in Fig. 9.
Figure 10: The realized gain of the antenna, which incorporates both the antenna gain and the effects of impedance matching, is a critical factor for achieving optimal performance.
The highest and flattest realized gain gives the best performance which provides some direction for improved future designs.
@@
IV-A2 Multiport Systems
The extension to multiport antennas or arrays is straightforward, as with dipole arrays within this section. The matrix formulation is also shown in (7), Appendices A and C.
The goal of this section is to show whether end-fire or broad-side pulsed radiation perform better in the setup of the four active dipoles discussed in Section III. The dipoles are spaced half a wavelength at $1\,\mathrm{GHz}$. Again, a band limited from above by frequency $3.8\,\mathrm{GHz}$ is considered.
The optimal radiated pulses of all configurations are compared in Fig. 12 where the magnitudes of the radiated Poynting vector peaks $S$ are given and normalized by radial distance $r$. The results show that broadside radiation performs better than end-fire.
Figure 12: Normalized Poynting vector of a single dipole antenna and array radiated in the end-fire and broadside direction.
Figure 13: Optimal excitation signals for end-fire and broadside pulsed radiation.
Considering the same input energy, the end-fire array increases the performance by more than a factor of four as compared to a single dipole. In the case of a broadside array, the performance boost reaches a factor of six. The difference may be due to the coupling of the field with the elements of the arrays which can be deduced from the optimal excitation waveforms in Fig. 13. Although the end-fire and broadside arrays share the same configuration, their characteristics, especially impedance ratios, differ since the dipoles are fed simultaneously. In this particular case, the broadside direction performs better than the end-fire orientation.
The excitations to all ports are almost identical in broadside operation. For the end-fire array, the delay equal to the distance between the dipoles multiplied by the speed of light is observed. The delays agree with the direction of end-fire radiation which points from dipole port 1 to dipole port 4. The magnitude of the excitation pulse is increased from dipole port 1 to dipole ports 2, 3, and 4, which indicates that the last element contributes the most.
The suggested approach can be applied to multi-dimensional arrays or any radiating structures with multiple ports and arbitrary configurations. Typically, computational demands grow as the number of elements increases. As demonstrated, expanding the number of independent ports results in expressing the excitation waveforms at each port through the basis functions described in Appendix B. The expansion coefficients are distinctively grouped into a vector that includes all dof.
IV-B Switching Antiferromagnetic Memory
The approach mentioned in this manuscript can be generalized in an arbitrary application where the input-output relation is known. In this section, the switching of an antiferromagnetic memory using a device illuminated by a pulsed laser is analyzed. The data are taken from [45]. The goal is to reach a peak electric field intensity high enough for memory switching.
@@ IV-B Switching Antiferromagnetic Memory
Figure 15: Incident plane waves at the position of the observation: (o) Pulse used in [45]. (a) Optimal excitation for the maximum peak. (b) Optimal excitation for the maximum peak with a constraint on the incident wave to concentrate $90\,\%$ of energy in the $2~{}\mathrm{ps}$ interval. (c) Optimal excitation for the maximum peak with a constraint in the target field to concentrate $90\,\%$ of energy in the $14~{}\mathrm{ps}$ interval. (d) Optimal excitation for the maximum peak with both constraints in (b) and (c).
The pulses are normalized to the maximum value of $10^{5}\,\mathrm{V/cm}$. The excited field magnitude in CuMnAs is provided in Fig. 16.
Figure 16: Magnitude of the electric field in CuMnAs for excitation pulses (o)–(d) in Fig. 15.
@@
One of the optimal pulses presented corresponds to the maximum magnitude of the electric field in CuMnAs in $t=0$ which is denoted as (a). The optimization problem is identical to (8) other than matrix $\mathbf{S}$ being replaced by $\mathbf{E}^{\mathrm{H}}\mathbf{E}$, where the electric field at the target point is represented as $\boldsymbol{E}_{\mathrm{CuMnAs}}(\omega)=\mathbf{E}(\omega)\mathbf{q}$ in the frequency domain, and $\mathbf{q}$ corresponds to the expansion of incident plan wave $a(\omega)$ into the basis functions of Appendix B, as done in (5). The relation is computed using CST which allows us to compute the electric field at a given point as a function of frequency considering unity excitation. Mapping the set of the given basis functions results in the desired matrix. The matrix representing the total energy of the incident plane wave is derived from (4) as demonstrated in (7), with the only variation being the inclusion of the free space impedance in the denominator.
@@
This evidence indicates the necessity of introducing constraints on the input and output waveforms. The incident wave is constrained to have $90\,\%$ energy in the interval $[-1.2,0.8]\,\mathrm{ps}$ taking into account the delay of $t_{\mathrm{d}}=0.2\,\mathrm{ps}$. The constraint is implemented using matrix $\mathbf{W}_{\mathrm{T}}$ of total available energy and energy available in the interval [2]
$W_{\mathrm{T}}=\dfrac{1}{4\pi^{2}Z_{0}}\int\limits_{-\infty}^{\infty}\int\limits_{-\infty}^{\infty}\dfrac{2\sin\left[\left(\omega_{1}-\omega_{2}\right)T\right]}{\omega_{1}-\omega_{2}}\mathrm{e}^{-\mathrm{j}\left(\omega_{1}-\omega_{2}\right)t_{\mathrm{d}}}\\ a^{*}\left(\omega_{2}\right)a\left(\omega_{1}\right)\mathrm{d}\omega_{2}\mathrm{d}\omega_{1},$ (9)
where $T=1\,\mathrm{ps}$ is the half-time interval and the asterisk represents a complex conjugation. In the frequency domain, the sinc function represents the process of time windowing. The matrix obtained is the outcome of the procedure illustrated in (7) regarding total input power. The corresponding QCQP reads
@@
where $\mathbf{U}$ and $\mathbf{U}_{\mathrm{T}}$ are matrices of the total squared electric field and the squared electric field in the given time interval for $T=7\,\mathrm{ps}$. The matrices are evaluated equivalently as $\mathbf{W}$ and $\mathbf{W}_{\mathrm{T}}$ except $t_{\mathrm{d}}$. The electric field at the target point as a function of frequency and unity excitation is taken from CST. The function is multiplied by its complex conjugate image and projected into the set of basis functions $\xi_{n}$ from Appendix B.
$\mathbf{U}=\dfrac{1}{4\pi}\int\limits_{-\infty}^{\infty}a^{*}\left(\omega\right)\boldsymbol{E}_{\mathrm{CuMnAs}}^{\mathrm{H}}(\omega)\boldsymbol{E}_{\mathrm{CuMnAs}}(\omega)a\left(\omega\right)\mathrm{d}\omega.$ (12)
Matrix $\mathbf{U}_{\mathrm{T}}$ differs only by introducing integration over two distinct angular frequencies $\omega_{1},\omega_{2}$ and the sinc function in the expression as in (9).
@@
V Conclusion
A methodology to determine optimal waveforms is presented in this paper with optimized parameters being defined as linear or quadratic terms. The evaluation requires a post-processing step based on convex optimization and the input data can be calculated in any EM simulator.
The tool can serve and provide quantitative results in many scenarios, including pulsed radiation or antiferromagnetic memory switching with pulse lasers, as discussed in this paper, as well as the creation of hotspots using near-field energy focusing in lossy materials using antenna arrays for applications such as hyperthermia therapy in medicine.

### v3

@@ IV Pulse Optimization Verification
The steps mentioned above are applied to an example of a single dipole antenna and instanenous far-field intensity $\boldsymbol{F}(\boldsymbol{d},t)$ in direction $\boldsymbol{d}$ orthogonal to the dipole orientation. Since this serves as verification, the redesign step S.6 is omitted, and codesign loop is presented in Section VII. The calculation is simultaneously performed in the Antenna Toolbox for MATLAB [31] (AToM) and CST. The dipole with a length of $L=150\,\mathrm{mm}$ and a width of $3\,\mathrm{mm}$ is oriented along the $z$ axis. The port is placed in the middle of the dipole and is connected to an ideal $50\,\mathrm{\Omega}$ feed line. Radiation in the spherical coordinate system is studied in the direction $\varphi=0$ and $\theta=\pi/2$. Far field consists of polarizations $\varphi$ and $\theta$. Azimuthal coordinate $\varphi$ references to angle from positive direction of $x$ axis in $xy$ plane and elevation coordinate $\theta$ references to angle from positive direction of $z$ axis.
@@
The antenna, including the port, is given with frequency range from $0\,\mathrm{GHz}$ to $3.8\,\mathrm{GHz}$ [25], and all steps visualized in Fig. 1 can be performed. Choosing AToM and CST as the EM solvers and computing the far-field along $\theta$ direction $F_{\theta}$ in elevation direction for unity excitation at all frequencies results in Fig. 2.
@@
Figure 2: Far field along $\theta$ direction $F_{\theta}$ of a dipole antenna with a length of $15\,\mathrm{mm}$ and width of $3\,\mathrm{mm}$ in directions $\varphi=0$ and $\theta=\pi/2$. The antenna is fed through a delta-gap source connected to the feed line, providing a voltage of $1\,\mathrm{V}$ across the entire frequency spectrum.
@@
Since the dipole is aligned along the $z$-axis, only the $\theta$ component of the electric field corresponds to radiation and far-field along direction $\phi$ is vanishingly small. The curves are in good agreement.
@@
In addition to the far field, an essential characteristic is total incident energy. This step accomplishes S.1.
@@
Since the frequency data from CST are dense due to the 1001 samples, only linear interpolation is used. Data resulting from AToM are sparse and rational fitting is used for interpolation [^8] Vector fitting [32] could be applied to enforce passivity and causality. according to S.2.
@@
Metrics such as far field and total input power are projected to the band-limited functions defined in Appendix C resulting in matrices as denoted in S.3,
$\boldsymbol{F}(\boldsymbol{d},t)=\mathbf{F}\mathbf{q}.$ (7)
The matrices allow for optimization S.4 which gives the optimal solution.
Thanks to basis function orthonormality, the identity matrix of the incident energy is used as the full rank positive definite matrix in the QCQP,
$\begin{split}\max\limits_{\mathbf{q}}\ &\mathbf{q}^{\mathrm{H}}\mathbf{F}^{\mathrm{H}}\mathbf{F}\mathbf{q},\\ \mathrm{s.t.}\ &\mathbf{q}^{\mathrm{H}}\mathbf{q}=W_{0}.\end{split}$ (8)
The optimal excitation is a complex conjugate of the far field [1] in Fig. 2. For presentation purposes, the total input energy was set to $W_{0}=10^{-10}\,\mathrm{J}$. Using S.5, the optimal waveform in the time domain is plotted in Fig.3.
@@
The last step is the computation of the radiated pulse, shown as the magnitude of far-field along $\theta$ direction $F_{\theta}$ in Fig. 4.
Figure 4: Far-field along $\theta$ direction $F_{\theta}$ optimal radiated waveform for $t_{0}=0\,\mathrm{s}$ at distance $r$ gained from iFT of the far field for AToM data and monitoring the far field during simulation in CST, where the optimal incident wave is applied as excitation.
@@
V Transmitting System
A representative example of a transmitting system is a single port antenna radiating in the target direction in the far field. The methodology can be used to find the one with the best performance, i.e., with the highest allowed intensity of the instantaneous field. As has been concluded in previous work on pulsed farfield radiation, the bowtie dipole antenna exhibits good performance among planar antennas studied for pulsed radiation in the direction orthogonal to the design plane [14, 16].
In this section, the following antennas are considered: a self-grounded bowtie antenna, a conical dipole antenna, a conical spiral antenna, and a logarithmic periodic antenna. The optimal waveform is determined to maximize the radiated peak power. Incident power waves are used as degrees of freedom, which allows us to consider matching (reflected power waves). This is essential for a realistic pulse response. This example aims to demonstrate the utility of the approach and compare the best potentially achievable performance of the proposed antennas, including their optimal excitation waveform.
@@ V Transmitting System
In [34], the conical spiral antenna is proposed for ground-penetrating radars, an UWB application. The conical spiral radiates a circularly polarized field compared to the other designs which operate in a linearly polarized regime. The antenna proposed in [34] is scaled to fit the circumscribing sphere mentioned above. The simulation is again performed in CST and the antenna is shown in Fig. 7.
@@
The periodic logarithmic antenna, recognized as an UWB antenna [35], is extensively utilized in electromagnetic emission and immunity testing [6]. Therefore, a wire dipole array is a potential candidate for pulsed radiation and is compared with other antennas. The array is designed to fit the circumscribing sphere and exhibits a decay of $0.9$ in element lengths and an opening angle of $15$ degrees. The strip width equals $0.017$ times the length of the longest element and the number of elements is 20. The paremeters are manualy tuned to achieve better performance. The antenna is simulated using AToM and shown in Fig. 8.
@@
The question remains as to which antenna performs better. A comparison of radiation intensity [36]
$U(t)=\dfrac{1}{Z_{0}}\left|\boldsymbol{F}(t)\right|^{2}$ (9)
is shown in Fig. 9.
Figure 9: Optimal radiation intensity $U$ at chosen distance $r$ from the compared antennas.
The logarithmic periodic antenna exceeds the other designs in performance. The performance of the self-grounded bowtie antenna is about two times lower and the other designs perform even worse. This finding could prompt efforts to develop improved designs. This may be achieved through redesign, involving iterative tuning of the logarithmic-periodic dipole wire antenna. An advanced topology optimization way is shown in Section VII using an automated approach on a planar antenna.
Regarding the origins of what constitutes good performance, the realized gain in the given direction accounts for the pulse’s sharpness and energy. The emerging differences in realized gain $G_{\mathrm{real}}$ of the antennas are depicted in Fig. 10 which justifies the performance differences presented in Fig. 9.
Figure 10: The realized gain $G_{\mathrm{real}}$ of the antenna, which incorporates both the antenna gain and the effects of impedance matching, is a critical factor for achieving optimal performance.
The highest and flattest realized gain gives the best performance which provides some direction for improved future designs. It is also noted that matching efficiency is a major part of the good realized gain of the logarithmic periodic antenna.
@@
VI Receiving Systems
Apart from transmission, the receiving performance is also covered by the approach. This shows that the method can also deal with the transmitting and receiving parts at the same time, but this is not shown in this manuscript.
In this section, focused on terahertz pulses, the switching of an antiferromagnetic memory using a device illuminated by a pulsed laser, is analyzed. The data are taken from [11]. The goal is to reach a peak electric field intensity high enough for memory switching.
@@ VI Receiving Systems
Figure 13: Time course of incident plane waves $E_{\mathrm{inc}}$ at the position of the observation: (o) Pulse used in [11]. (a) Optimal excitation for the maximum peak. (b) Optimal excitation for the maximum peak with a constraint on the incident wave to concentrate $90\,\%$ of energy in the $2~{}\mathrm{ps}$ interval. (c) Optimal excitation for the maximum peak with a constraint in the target field to concentrate $90\,\%$ of energy in the $14~{}\mathrm{ps}$ interval. (d) Optimal excitation for the maximum peak with both constraints in (b) and (c).
The pulses are normalized to the maximum value of $10^{5}\,\mathrm{V/cm}$. The excited field magnitude in CuMnAs $E_{\mathrm{CuMnAs}}$ is provided in Fig. 14.
Figure 14: Magnitude of the electric field in CuMnAs $E_{\mathrm{CuMnAs}}$ for excitation pulses (o)–(d) in Fig. 13.
@@
One of the optimal pulses presented corresponds to the maximum magnitude of the electric field in CuMnAs in $t=0$ which is denoted as (a). The optimization problem is identical to (8), where $\mathbf{q}$ corresponds to the expansion of incident plane wave $a(\omega)$ into the basis functions of Appendix C, as done in (3). The relation is computed using CST which allows us to compute the electric field at a given point as a function of frequency considering unity excitation. Mapping the set of the given basis functions results in the desired matrix. The matrix representing the total energy of the incident plane wave is derived in Appendix B.
@@
This evidence indicates the necessity of introducing constraints on the input and output waveforms. The incident wave is constrained to have $90\,\%$ energy in the interval $[-1.2,0.8]\,\mathrm{ps}$ taking into account the delay of $t_{\mathrm{d}}=0.2\,\mathrm{ps}$. The constraint is implemented using matrix derived from operator given in Appendix B. The corresponding QCQP reads
@@
where $\mathbf{U}$ and $\mathbf{U}_{\mathrm{T}}$ are matrices of the total squared electric field and the squared electric field in the given time interval $T=7\,\mathrm{ps}$ similarly derived as delivered energy matrices in Appendix B.
@@
VII Optimization of the System’s Topology
Since the excitation waveform optimization is fast and robust, it can be combined with the topology optimization. The target application is antenna and far-field radiation. The antenna is a multiport system, and this example, as a by-product, proves its applicability to systems with multiple inputs [^9] Multiple output systems within the objective can be treated by multiobjective optimization [30], whose solutions represent Pareto optimal points..
The initial example consists of an array of two parallel and identical dipoles with the same dimensions as in Section IV. The frequency band also remains the same. The goal is to maximize instantaneous radiation in the end-fire direction.
The simplicity of the design enables rapid assessment of the impulse response and facilitates the transfer of data between iterations without requiring re-evaluation in a full-wave solver, which is beneficial when integrating system and waveform optimization. Application to larger structures with multiple ports does not bring any complications except for computation time.
By performing a parametric sweep, the optimal spacing between the dipoles is determined. Different configurations are evaluated:
  • Both dipoles are driven.
  • One dipole is driven, the other is shorted and acts as a reflector.
  • One dipole is driven, the other is shorted and acts as a director.
  • One dipole is driven and the reflector is, in its geometrical center, loaded by an optimized reactive loading.
The maximum instantaneous radiation intensity achieved by these setups is shown in Fig. 15 as a function of the distance between the dipoles, which is normalized to twice the dipole length, corresponding to the wavelength for which the dipole is designed.
Figure 15: Comparison of maximum instantaneous radiation intensity $U$ for an end-fire two-dipole array for different distances between the dipoles normalized by the length of the dipoles. Performance is benchmarked against a single dipole. Considered configurations include both elements being driven, one element being driven with either a shorted director or reflector, and a reflector with an optimized load or a reflector setup with optimized topology. In the top panel, the solid lines show performance without considering impedance mismatch, reactive loading, or shape optimization, while, in the bottom panel, the dashed lines account for practical matching to an ideal $50\,\Omega$ feeding line, providing a more realistic assessment. The dotted line in the bottom panel represents a dipole reflector with an optimized reactive load, and the markers in the bottom panel represent reflector setups with optimized shapes (a)-(d) in Fig. 16. Both optimizations in the bottom panel consider connection to ideal $50\,\Omega$ feed line.
[…]
Figure 16: Topologically optimized antenna shapes with ports highlighted in green. Instantaneous radiation intensity in the north direction is optimized (reflector setup):
Shape (a) Best performance among the presented designs.
Shape (b) Elements of the reflector closest to the driven element.
Shape (c) Good trade-off between the shape regularity and performance.
Shape (d) The simplest and the most compact design.
Figure 15 shows two panels:
  • In the top panel, the solid lines with better performance omit the matching (assume ideal voltage generators at all frequencies). The dof are delta-gap voltages. In that case, the instantaneous far-field intensity is normalized to the total dissipated energy due to ohmic losses in the elements and radiation.
  • In the bottom panel, the dashed lines with lower performance consider the reflection of waves at the ports i.e., matching. The dof are incident power waves in the $50\,\Omega$ feeding lines. There, the instantaneous far-field intensity is normalized to the total incident energy.
As expected, the best performance is achieved when both dipoles are fed independently. A similar trend is followed by a single fed dipole with a shorted reflector; for larger distances, the impact of the reflector is negligible, and the performance is identical to that of a dipole alone. The director decreases performance as compared to a single dipole.
Continuing in the co-design, the performance of the setup with reflector can be improved by loading the reflector with a shorted ideal $50\,\Omega$ line of an optimized length $\ell$ in range from $0\,\mathrm{m}$ to $2L$ (reactive loading). The optimization of reactive loading is only considered in the realistic scenario, where reflection caused by mismatch is taken into account, and the optimized dof is the incident power wave. The enhancement is significant at small separations between the driven dipole and the reflector. The optimized design performance is plotted as a dotted line in Fig. 15 and can be compared to the dashed line for the shorted reflector. The best possible performance is $36.2\,\mathrm{mW}$ which corresponds to $d/(2L)=0.11$ and line length $\ell/(2L)=0.20$. In the same way, the load on the director was also optimized, but the performance did not exceed that of a single dipole, so the result is not depicted.
The ultimate solution is co-design with the topology optimization [37, 38], which is implemented for the reflector setup. Performances are shown as markers in Fig. 15 for designs (a) – (d) from Fig. 16, respectively. The antennas possess east-west mirror symmetry with respect to the north target direction. Designs are ranked according to their performance, inversely related to their regularity [39]. Performance enhancement is significant and underscores the effectiveness of the co-design methodology. This approach effectively addresses mismatch issues, improves gain, and exceeds the capabilities of reactive loading in terms of functionality.
VIII Conclusion